OpenStudy (anonymous):

Which equation models this problem? Colin divided his baseball card collection into 6 equal piles. He put the cards from 5 of the piles into albums. How many cards are in his collection if he put 140 baseball cards into albums?

OpenStudy (anonymous):

A. \[\frac{ 140 }{ 6 }b=5\] B. \[\frac{ 140 }{ 5 }b=\epsilon \] C. \[\frac{ 5 }{ 6 }b = 140\] D. \[\frac{ 6 }{ 5 }b=140\] @kropot72 @iambatman

OpenStudy (anonymous):

@kobeni-chan

OpenStudy (kobeni-chan):

B.

OpenStudy (kaydra14):

i think the answer is B